# Scalable Kernel Logistic Regression with Nyström Approximation: Theoretical Analysis and Application to Discrete Choice Modelling

### v1

@@ 1 Introduction
KLR has two crucial advantages in the area of model output derivation. Firstly, it operates with latent functions which take on the role of utility functions in traditional DCM, but with the notable advantage of not requiring a pre-specification of the utility function. This allows for the automatic capture of non-linear effects, leading to greater flexibility in modelling. Secondly, KLR excels at properly estimating choice probabilities and their derivatives with respect to the attributes, providing essential information for behavioural analysis of the various models (Martín-Baos et al., 2021). These advantages collectively enhance the suitability and performance of KLR in transportation mode choice analysis.
@@ 2 Related work
To increase the efficiency of the method, common approaches involve employing approximation techniques that yield a low-rank representation of the kernel matrix. These methods include subset of regressors (Zhang et al., 2011), the Nyström method (Lei et al., 2019), subset of data (Zhu and Hastie, 2005), and incomplete Choleski factorisation (Cawley and Talbot, 2005; López García et al., 2015), among others. The focus of this study is on the Nyström method, a technique that has achieved excellent results in the approximation of the kernel matrix (Sun et al., 2015), making it a commonly-used approach in other domains.
@@
The Nyström method’s efficacy has been demonstrated in Kernel Regression applications, as evidenced by works such as Alaoui and Mahoney (2015) and Zhang et al. (2022). To enhance the accuracy of the Nyström approximation, several improved variants have been developed, including ensemble Nyström (Sun et al., 2015), modified Nyström (De Bonis and Laurita, 2016), and SS-Nyström (Wang et al., 2014). Another approach involves applying sampling analysis methods to identify a subset of landmark points that capture the majority of the dataset’s information. These are then used for the Nyström computation. A detailed exploration of these sampling methods can be found in Sun et al. (2015).
@@
The work of Gittens and Mahoney (2016) highlights that the performance of various Nyström approximations depends on intricate factors present in the data structure, such as spectral decay, leverage score structure, eigenvalue gaps in relevant parts of the spectrum, etc. Zhang et al. (2022) examined a non-uniform Nyström approximation for sparse kernel regression and introduce two non-uniform Nyström methods with theoretical guarantees for sparse kernel regression. While theoretical contributions in prior works can offer recommendations on the suitability of different methods, the most appropriate technique for a particular situation is ultimately determined by the specific application. Regarding Nyström’s application to KLR, Lei et al. (2019) proposed combining the basic Nyström method with a fast iterative algorithm, similar to sequential minimal optimisation, to solve binary classification problems. However, the studies conducted by these authors addressed simpler problems compared to those encountered in this research. Specifically, they exclusively dealt with binary classification problems, and the datasets they considered contained fewer than $4,000$ samples.
@@ 3 Multinomial Kernel Logistic Regression
The foundation of KLR lies in constructing several latent functions, which are analogous to the systematic utility functions found in RUMs. However, unlike RUM, KLR treats these latent functions as black boxes, where the modeller does not need to explicitly state the relationship between the feature vector and the utilities. Rather, KLR operates within the framework of RKHS, seeking these latent functions $f_{i}:\mathcal{X}\mapsto\mathbb{R}$ for $i=1,\ldots,I$ within this function space. The RKHS is univocally generated by the positive-definite real-valued kernel on $k:{\cal X}\times{\cal X}\mapsto\mathbb{R}$, and its associated RKHS space is denoted by ${\mathcal{H}}_{k}$. The family of functions $\{k(\mathbf{x},\mathbf{x}^{\prime})\}_{\mathbf{x}^{\prime}\in\mathcal{X}}$ constitutes a basis for this vector space. Thus, any element from ${\mathcal{H}}_{k}$ can be represented as a linear combination of basis elements, in particular for $f_{i}\in{\mathcal{H}}_{k}$.
The utilisation of RKHS ensures a flexible and powerful approach to modelling the complex relationships between the feature vectors and latent functions, making KLR a highly effective classification tool in various applications. Knowing these latent functions, the classification criterion used is:
$f(\mathbf{x})=\underset{i\in\{1,\ldots,I\}}{\hbox{\rm arg maximise }}f_{i}(\mathbf{x}).$ (1)
where arg maximise denotes the input point $i$ at which the function output values $f_{i}(\mathbf{x})$ is maximised.
Let $\{\mathbf{x}_{n},y_{n}\}_{n=1}^{N}$ be $N$ pairs of points in ${\cal X}\times{\cal Y}$, where ${\cal X}\subseteq{\mathbb{R}}^{M}$ is the input space and $\cal Y\subseteq\mathbb{R}$ is the response space, which, in our context, serves as a label space. The kernel-based learning problem can be stated as the following optimisation problem:
$\underset{f\in{\cal H}_{k}}{\hbox{Minimise }}\frac{1}{N}\sum_{n=1}^{N}\ell(y_{n},f(\mathbf{x}_{n}))+R(f),$ (2)
where $\ell:{\cal Y}\times{\cal Y}\mapsto\mathbb{R}$ is a loss function that quantifies the discrepancies between the predicted, $f(\mathbf{x}_{n})$, and the observed labels, $y_{n}$, and $R(f)$ is a regularisation term to enhance the generalisation capacity of the model.
@@
if and only if there exists a non-decreasing function $h:[0,\infty)\mapsto\mathbb{R}$ for which $R(f)=h(\|f\|_{{\cal H}_{k}})$. As a straightforward consequence, the problem represented in Equation (2) is reduced to an finite-dimensional optimisation problem where the goal is to find a matrix $\boldsymbol{\alpha}_{\cdot\cdot}=(\alpha_{ni})\in\mathbb{R}^{N\times I}$ that solves:
$\underset{\boldsymbol{\alpha_{\cdot\cdot}}\in\mathbb{R}^{N\times I}}{\hbox{Minimise }}\frac{1}{N}\sum_{n=1}^{N}\ell(y_{n},f(\mathbf{x}_{n}))+R(f),$ (4)
One of the most commonly used regularisation term is the Ridge penalisation, often referred to as Tikhonov regularisation. This regularisation term includes a convex and differentiable quadratic component in the objective function, which tends to smooth out the model’s parameters and reduce their variance, thereby helping to prevent overfitting and enhancing the generalisation capacity of the model. This, in turn, allows the model to perform better on unseen data. Tikhonov regularisation is defined as the sum of the norm of the latent functions in RKHS:
$R(f)=\dfrac{\lambda}{2}\sum_{i=1}^{I}\|f_{i}(\mathbf{x})\|^{2}_{{\cal H}_{K}}=\dfrac{\lambda}{2}\sum_{i=1}^{I}\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}k(\mathbf{x}_{n},\mathbf{x}_{n^{\prime}})\alpha_{ni}\alpha_{n^{\prime}i}=\dfrac{\lambda}{2}\sum_{i=1}^{I}\boldsymbol{\alpha}_{\cdot i}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\cdot i},$ (5)
where $\mathbf{K}$ is the so-called kernel or Gram matrix, defined as follows: $\mathbf{K}_{n\,n^{\prime}}=k(\mathbf{x}_{n},\mathbf{x}_{n^{\prime}})\hbox{ for}n,n^{\prime}=1,\ldots,N;$ and $\boldsymbol{\alpha}_{\cdot i}$ is the $i-$th column of the matrix $\boldsymbol{\alpha}_{\cdot\cdot}$, i.e. $\boldsymbol{\alpha}_{\cdot\cdot}=(\boldsymbol{\alpha}_{\cdot 1},\ldots,\boldsymbol{\alpha}_{\cdot I})\in\mathbb{R}^{N\times I}$. As $k(\cdot,\cdot)$ is a kernel function, the Gram matrix is guaranteed to be symmetric and positive definite. Additionally, the regularisation parameter $\lambda$ controls the trade-off between the model’s goodness of fit and its complexity. In order to facilitate the later computation of the gradient, the regularisation parameter is divided by $2$ in this expression.
@@
Hence, the RNLL problem used for training the KLR can be formulated as follows:
$\underset{\boldsymbol{\alpha}_{\cdot\cdot}}{\hbox{Minimise }}{\cal L}(\boldsymbol{\alpha}_{\cdot\cdot})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{I}y_{ni}\log p_{ni}(\boldsymbol{\alpha}_{\cdot\cdot})+\frac{\lambda}{2}\sum_{i=1}^{I}\boldsymbol{\alpha}_{\cdot i}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\cdot i},$ (7)
where the matrix $\mathbf{y}_{\cdot\cdot}$ represents whether individual $n$ has selected alternative $i$ (i.e. it is the one-hot encoding of the label assigned to $n$, $y_{ni}=1$ if $y_{n}=i$ and $y_{ni}=0$ otherwise).
@@ 3.1 Preliminaries
$\|\mathbf{A}\|_{\infty}:=\underset{\|\mathbf{a}\|_{2}=1}{\max}\frac{\|\mathbf{A}\mathbf{a}\|_{2}}{\|\mathbf{a}\|_{2}}=\sqrt{(\sigma_{1}(\mathbf{A}^{\top}\mathbf{A})}$
@@ 3.2 Analysis of the solution set of the KLR
where $\boldsymbol{\eta}_{\cdot\cdot}\in L(\mathbf{K})^{\perp}$, *$\boldsymbol{\aleph}^{*}_{\cdot\cdot}\in L(\mathbf{K})$, and $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ is the unique solution to the problem (7) restricted to the subspace $L(\mathbf{K})$:
@@
$\mathbf{K}\left[N\lambda\boldsymbol{\alpha}^{*}_{\cdot i}-\mathbf{p}_{\cdot i}(\boldsymbol{\alpha}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots I.$ (15)
@@ 4.1 Matrix sketching and the Nyström Method
The Nyström method exploits the assumption that certain individuals (or observations) in the kernel matrix contain redundant information (i.e. rank$(\mathbf{K})<N$), and thus it is convenient to avoid duplicating this information. Then, the goal is to select the most representative columns with the so-called sketching matrix. Formally, given the kernel matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$, it is called $\mathbf{C}=\mathbf{K}\mathbf{P}\in\mathbb{R}^{N\times C}$ (typically $C\ll N$) as a sketch of $\mathbf{K}$ and $\mathbf{P}\in\mathbf{R}^{N\times C}$ is a sketching matrix. The columns of $\mathbf{C}$ typically comprise a rescaled subset of the columns of $\mathbf{K}$; the former type of sketching is called column selection111 An alternative approach, distinct from the one examined in the paper, involves considering random linear combinations of the columns of matrix $\mathbf{K}$, commonly referred to as a random projection. .
@@
where $\mathbf{W}=\mathbf{P}^{\top}\mathbf{K}\mathbf{P}$ and ${}^{\dagger}$ denotes the Moore-Penrose pseudoinverse. The first thing to note about Equation (16) is that the columns of the matrix $\hat{\mathbf{K}}$ are a linear combination of the columns of $\mathbf{C}$, so that $L(\mathbf{\widehat{K}})\subset L(\mathbf{C})$. Since both vector spaces have the same dimension, $L(\widehat{\mathbf{K}})=L(\mathbf{C})$.
@@
The Nyström approximation $\widehat{\mathbf{K}}$ is the best possible $C-$ approximation to $\mathbf{K}$ given the information $\mathbf{C}$.
In more detail, for all symmetric matrices $\widehat{\mathbf{K}}$ spanned by the columns of $\mathbf{C}$ with a positive semi-definite residual matrix $\mathbf{K}-\widehat{\mathbf{K}}$, the Nyström approximation has the smallest error as measured by either any unitary invariant norm, such as the Schatten $-p$ norm.

### v3

@@ 1 Introduction
KLR has two crucial advantages in the area of model output derivation. Firstly, it operates with latent functions which take on the role of utility functions in traditional DCM, but with the notable advantage of not requiring a pre-specification of the utility function. This allows for the automatic capture of non-linear effects, leading to greater flexibility in modelling (Martín-Baos et al., 2020). Secondly, KLR excels at properly estimating choice probabilities and their derivatives with respect to the attributes, providing essential information for behavioural analysis of the various models (Martín-Baos et al., 2021). These advantages collectively enhance the suitability and performance of KLR in transportation mode choice analysis.
@@ 2 Related work
To increase the efficiency of the method, common approaches involve employing approximation techniques that yield a low-rank representation of the kernel matrix. These methods include subset of regressors (Zhang et al., 2011), the Nyström method (Lei et al., 2019), subset of data (Zhu and Hastie, 2005), and incomplete Choleski factorisation (Cawley and Talbot, 2005; López García et al., 2015), among others.
Nyström’s methods have been successfully applied to large-scale machine learning techniques such as Kernel Regression applications (Alaoui and Mahoney, 2015; Zhang et al., 2022), numerical solutions of an integral equations system (Allouch et al., 2021), among other domains. This large number of applications motivates a research effort per se to obtain error bounds for low-rank approximation of kernel matrices $\mathbf{K}$. Fundamental theoretical results on the error bounds of the Nyström approximation are presented in Zhang et al. (2008); Kumar et al. (2012); Sun et al. (2015); Gittens and Mahoney (2016). These studies highlight that the performance of various Nyström approximations depends on intricate factors present in the data structure, such as spectral decay, leverage score structure, eigenvalue gaps in significant parts of the spectrum, etc.
@@
The analysis of these bounds allows for the development of methods to enhance the accuracy of the Nyström approximation, and therefore several improved variants have been developed, including ensemble Nyström (Sun et al., 2015), modified Nyström (De Bonis and Laurita, 2016), and SS-Nyström (Wang et al., 2014). Another approach involves applying sampling analysis methods in order to identify a subset of landmark points that capture the majority of the dataset’s information. A detailed exploration of these sampling methods can be found in Sun et al. (2015).
In the application of Nyström methods to machine learning, it is desired to know how the error of the approximation affects the performance of the algorithm itself. These studies are usually performed in two stages, the first involves bounding the error of the learnt parameters as a function of the Nyström approximation error, and the second involves bounding the learning performance based on the errors of the resulting parameters. In Cortes et al. (2010) the error bounds for SVM, Kernel Ridge Regression (KRR), and graph Laplacian-based regularisation algorithms were studied. Zhang et al. (2022) examined a non-uniform Nyström approximation for Sparse Kernel Regression (SKR), and introduced two non-uniform Nyström methods with theoretical guarantees for SKR. Wang et al. (2019) analysed the application of the so-called rank-restricted Nyström approximation to kernel $k-$means clustering and showed that it satisfied a $1+\varepsilon$ approximation ratio in terms of the kernel $k-$means cost function, related to the guarantee provided by the same algorithm without using the Nyström method.
To the best of our knowledge, no bounds have been established for the Nyström KLR and this paper analyses the problem. In this study, theoretical bounds have been established for the estimated parameters and it has been numerically observed that there is a correlation between these bounds and the usual metrics used in the validation of the KLR.
@@
Regarding Nyström’s application to KLR, Lei et al. (2019) proposed combining the standard Nyström method with a fast iterative algorithm, similar to sequential minimal optimisation, to solve binary classification problems. However, the studies conducted by these authors addressed simpler problems compared to those encountered in this research. Specifically, they exclusively dealt with binary classification problems, and the datasets they considered contained fewer than $4,000$ samples.
@@ 3 Multinomial Kernel Logistic Regression
The foundation of KLR lies in constructing several latent functions, which are analogous to the systematic utility functions found in RUMs. However, unlike RUM, KLR treats these latent functions as black boxes, where the modeller does not need to explicitly state the relationship between the feature vector and the utilities. Rather, KLR operates within the framework of RKHS, seeking these latent functions, which are denoted by $f_{i}:\mathcal{X}\mapsto\mathbb{R}$ for each alternative $i=1,\ldots,I$, within this function space. The RKHS is univocally generated by the positive-definite real-valued kernel on $k:{\cal X}\times{\cal X}\mapsto\mathbb{R}$, and its associated RKHS space is denoted by ${\mathcal{H}}_{k}$. The family of functions $\{k(\mathbf{x},\mathbf{x}^{\prime})\}_{\mathbf{x}^{\prime}\in\mathcal{X}}$ constitutes a basis for this vector space. Thus, any element from ${\mathcal{H}}_{k}$ can be represented as a linear combination of basis elements, in particular for $f_{i}\in{\mathcal{H}}_{k}$.
The utilisation of RKHS ensures a flexible and powerful approach to modelling the complex relationships between the feature vectors and latent functions, making KLR a highly effective classification tool in various applications. Given these latent functions, the classification criterion used for a given sample $\mathbf{x}$ is:
$\widetilde{f}(\mathbf{x})=\underset{i\in\{1,\ldots,I\}}{\hbox{\rm arg maximise}}f_{i}(\mathbf{x}).$ (1)
where arg maximise denotes the alternative $i$ at which the function output values $f_{i}(\mathbf{x})$ is maximised. The latent functions $f_{i}(\mathbf{x})$ serve as an index of how well the sample $\mathbf{x}$ fits alternative $i$, and the classifier $\widetilde{f}(\mathbf{x})$ assigns the alternative for which this index is highest.
Let $\{\mathbf{x}_{n},y_{n}\}_{n=1}^{N}$ be $N$ pairs of points in ${\cal X}\times{\cal Y}$, where ${\cal X}\subseteq{\mathbb{R}}^{M}$ is the input space and $\cal Y$ is the response space, which, in our context, serves as a label space. The kernel-based learning problem can be stated as the following optimisation problem:
$\underset{\widetilde{f}}{\hbox{Minimise }}\frac{1}{N}\sum_{n=1}^{N}\ell\left(y_{n},\widetilde{f}(\mathbf{x}_{n})\right)+R(\widetilde{f}),$ (2)
where $\ell:{\cal Y}\times{\cal Y}\mapsto\mathbb{R}$ is a loss function that quantifies the discrepancies between the predicted, $\widetilde{f}(\mathbf{x}_{n})$, and the observed labels, $y_{n}$, and $R(\widetilde{f})$ is a regularisation term to enhance the generalisation capacity of the model.
@@
if and only if there exists a non-decreasing function $h:[0,\infty)\mapsto\mathbb{R}$ for which $R(\widetilde{f})=h\left(\|\widetilde{f}\|\right)$. It can be seen how $f_{i}(\mathbf{x})$ depends on the parameter vector $\boldsymbol{\alpha}_{\cdot i}$ and in turn, the classifier $\widetilde{f}(\mathbf{x})$ depends on all latent functions. It will therefore also depend on this set of parameter vectors which will be gathered in the form of a matrix by writing $\boldsymbol{\alpha}_{\cdot\cdot}=(\alpha_{ni})\in\mathbb{R}^{N\times I}$. As a straightforward consequence, the problem represented in Equation (2) is reduced to an finite-dimensional optimisation problem where the goal is to find a matrix $\boldsymbol{\alpha}_{\cdot\cdot}$ that solves:
$\underset{\boldsymbol{\alpha_{\cdot\cdot}}\in\mathbb{R}^{N\times I}}{\hbox{Minimise }}\frac{1}{N}\sum_{n=1}^{N}\ell(y_{n},\widetilde{f}(\mathbf{x}_{n}))+R(\widetilde{f}),$ (4)
One of the most commonly used regularisation terms, $R(\widetilde{f})$, is the Ridge penalisation, often referred to as Tikhonov regularisation. This regularisation term includes a convex and differentiable quadratic component in the objective function, $\dfrac{\lambda}{2}\sum_{i=1}^{I}\boldsymbol{\alpha}_{\cdot i}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\cdot i}$, where $\mathbf{K}$ is the so-called kernel or Gram matrix, defined as follows: $\mathbf{K}_{n\,n^{\prime}}=k(\mathbf{x}_{n},\mathbf{x}_{n^{\prime}})\hbox{ for}n,n^{\prime}=1,\ldots,N;$ and $\boldsymbol{\alpha}_{\cdot i}$ is the $i-$th column of the matrix $\boldsymbol{\alpha}_{\cdot\cdot}$, i.e. $\boldsymbol{\alpha}_{\cdot\cdot}=(\boldsymbol{\alpha}_{\cdot 1},\ldots,\boldsymbol{\alpha}_{\cdot I})\in\mathbb{R}^{N\times I}$.
Since $k(\cdot,\cdot)$ is a kernel function, the Gram matrix $\mathbf{K}$ is guaranteed to be symmetric and positive semi-definite, ensuring the convexity of $R(\widetilde{f})$. This regularisation term tends to smooth out the model’s parameters and reduce their variance, thereby helping to prevent overfitting and enhancing the generalisation capacity of the model (Hastie et al., 2001). This, in turn, allows the model to perform better on unseen data.
Tikhonov regularisation is defined as the sum of the norm of the latent functions in RKHS:
$R(\widetilde{f})=\dfrac{\lambda}{2}\sum_{i=1}^{I}\|f_{i}(\mathbf{x})\|^{2}_{{\cal H}_{K}}=\dfrac{\lambda}{2}\sum_{i=1}^{I}\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}k(\mathbf{x}_{n},\mathbf{x}_{n^{\prime}})\alpha_{ni}\alpha_{n^{\prime}i}=\dfrac{\lambda}{2}\sum_{i=1}^{I}\boldsymbol{\alpha}_{\cdot i}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\cdot i},$ (5)
where the regularisation parameter $\lambda$ controls the trade-off between the model’s goodness of fit and its complexity. In order to facilitate the later computation of the gradient, the regularisation parameter is divided by $2$ in this expression.
@@
Henceforth, the matrix $\mathbf{y}_{\cdot\cdot}$ is defined to indicate whether individual $n$ has selected alternative $i$ using the one-hot encoding of $y_{n}$. Specifically, $y_{ni}=1$ if $y_{n}=i$ and $y_{ni}=0$ otherwise. Consequently, the RNLL problem used for training the KLR can be formulated as follows:
$\underset{\boldsymbol{\alpha}_{\cdot\cdot}}{\hbox{Minimise }}{\cal L}(\boldsymbol{\alpha}_{\cdot\cdot})=\underset{\boldsymbol{\alpha}_{\cdot\cdot}}{\hbox{Minimise }}\left(-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{I}y_{ni}\log p_{ni}(\boldsymbol{\alpha}_{\cdot\cdot})+\frac{\lambda}{2}\sum_{i=1}^{I}\boldsymbol{\alpha}_{\cdot i}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\cdot i}\right).$ (7)
@@ 3.1 Preliminaries
$\|\mathbf{A}\|_{\infty}:=\underset{\|\mathbf{a}\|_{2}=1}{\max}\frac{\|\mathbf{A}\mathbf{a}\|_{2}}{\|\mathbf{a}\|_{2}}=\sqrt{\sigma_{1}(\mathbf{A}^{\top}\mathbf{A})}$
@@ 3.2 Analysis of the solution set of the KLR
where $\boldsymbol{\eta}_{\cdot\cdot}\in L(\mathbf{K})^{\perp}$, $\boldsymbol{\aleph}^{*}_{\cdot\cdot}\in L(\mathbf{K})$, and $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ is the unique solution to the problem (7) restricted to the subspace $L(\mathbf{K})$:
@@
$\mathbf{K}\left[N\lambda\boldsymbol{\alpha}^{*}_{\cdot i}+\mathbf{p}_{\cdot i}(\boldsymbol{\alpha}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots I.$ (15)
@@ 4.1 Matrix sketching and the Nyström Method
The Nyström method exploits the assumption that certain individuals (or observations) in the kernel matrix contain redundant information (i.e. rank$(\mathbf{K})<N$), and thus it is convenient to avoid duplicating this information. Then, the goal is to select the most representative columns with the so-called sketching matrix. Formally, given the kernel matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$, it is called $\mathbf{C}=\mathbf{K}\mathbf{P}\in\mathbb{R}^{N\times C}$ (typically $C\ll N$) as a sketch of $\mathbf{K}$ and $\mathbf{P}\in\mathbf{R}^{N\times C}$ is a sketching matrix. The columns of $\mathbf{C}$ typically comprise a rescaled subset of the columns of $\mathbf{K}$; the former type of sketching is called column selection. An alternative approach involves considering random linear combinations of the columns of matrix $\mathbf{K}$, commonly referred to as a random projection.
@@
where $\mathbf{W}=\mathbf{P}^{\top}\mathbf{K}\mathbf{P}$ and denotes the Moore-Penrose pseudoinverse. The first thing to note about Equation (16) is that the columns of the matrix $\hat{\mathbf{K}}$ are a linear combination of the columns of $\mathbf{C}$, so that $L(\mathbf{\widehat{K}})\subseteq L(\mathbf{C})$. Furthermore, $\hbox{rank}(\mathbf{\widehat{K}})=\hbox{rank}(\mathbf{C})$ [^1] First calculate a bound for the rank of the pseudoinverse of $\mathbf{W}$, $\hbox{rank}(\mathbf{W}^{\dagger})=\hbox{rank}(\mathbf{W})=\hbox{rank}((\mathbf{K}^{1/2}\mathbf{P})^{\top}(\mathbf{K}^{1/2}\mathbf{P}))\geq\hbox{rank}((\mathbf{K}^{1/2}\mathbf{P})^{\top}\mathbf{K}(\mathbf{K}^{1/2}\mathbf{P}))=\hbox{rank}(\mathbf{C}^{\top}\mathbf{C})=\hbox{rank}(\mathbf{C}).$ It follows that: $\hbox{rank}(\widehat{\mathbf{K}})=\hbox{rank}(\mathbf{D}^{\top}\mathbf{D})=\hbox{rank}(\mathbf{D}^{\top})$ where $\mathbf{D}^{\top}=\mathbf{C}(\mathbf{W}^{\dagger})^{1/2}$. On the other hand, $\mathbf{W}^{\dagger}$ is symmetric and positive semidefinite, therefore $\hbox{rank}(\mathbf{W}^{\dagger})=\hbox{rank}((\mathbf{W}^{\dagger})^{1/2}(\mathbf{W}^{\dagger})^{1/2})=\hbox{rank}((\mathbf{W}^{\dagger})^{1/2}).$ It then follows that $\hbox{rank}((\mathbf{W}^{\dagger})^{1/2})\geq\hbox{rank}(\mathbf{C})$. If it is observed that $\mathbf{D}^{\top}$ is formed by more than rank$(\mathbf{C})$ independent linear combinations of all the columns of $\mathbf{C}$, it must therefore have a rank greater than or equal to rank$(\mathbf{C})$. Since the maximum rank it can have is $\hbox{rank}(\mathbf{C})$ it follows that rank$(\widehat{\mathbf{K}})=\hbox{rank}(\mathbf{C})$. and from there it follows that $L(\widehat{\mathbf{K}})=L(\mathbf{C}).$
@@
[…]